OpenStudy (anonymous):

How do I solve the question? If a/4=b/7, what is the value of a/b?

OpenStudy (anonymous):

Thanks alot!

OpenStudy (anonymous):

\[\frac{a}{4}=\frac{b}{7}\] \[7a=4b\] \[a=\frac{4b}{7}\] \[\frac{a}{b}=\frac{7}{4}\]

OpenStudy (anonymous):

what @satellite73 ?

OpenStudy (amistre64):

the problem with people just posting silly crap from the first thing they find on the internet is that usually they have no idea what they are talking about :/ the first poster didnt even answer your question but most likely copy and pasted it from: http://www.enotes.com/math/q-and-a/4-b-7-calculate-value-a-b-a-b-190067

OpenStudy (anonymous):

Oh okay...